# The Primordial Inflation Explorer (PIXIE): Mission design and science goals

### v1

@@ 3.3 Particle Physics
The neutrino is the only Standard Model particle whose mass is unknown. Measurements of E-mode polarization enable detection of the neutrino mass. Neutrinos affect the growth of structure: massive neutrinos cannot cluster on physical scales smaller than their free-streaming length, suppressing the growth of structure on these scales. Gravitational lensing from CMB and optical surveys determines the amplitude of current structure in our universe, but determining the growth of structure requires knowledge of the primordial amplitude as well. Primary CMB anisotropy only determines the primordial amplitude within $\sim$1uncertainty due to scattering at reionization. PIXIE’s nearly cosmic-variance limited measurement of the reionization optical depth (§$\S$2.2) would remove this limiting uncertainty to enable a $4\sigma$ determination of neutrino mass [80].
Dark matter (DM) provides a more exotic test of particle physics. The long-favored weakly interacting massive particles (WIMP) scenario is seeing increased pressure from direct detection experiments, suggesting instead more exotic scenarios such as axionic DM [81, 82] or PBHs [83]. PIXIE’s spectral distortion limits constrain electromagnetic interactions between dark matter and standard-model particles from DM decay annihilation and scattering [84]. Spectral distortions from DM decay depend on the lifetime and abundance of the (excited) DM state. The temperature/velocity- dependence of the DM annihilation cross-section strongly affects annihilation signals, with spectral distortions being particularly valuable for p-wave annihilation [85]. Since the dark matter number density scales as 1/mass, spectral distortions from DM annihilation or scattering are sensitive to low masses. PIXIE 95% confidence limits $|\mu|<6\times 10^{-7}$ from the baseline mission restrict Majorana annihilations to mass $m_{\chi}>{\sim}300$ keV [86].
@@ 3.4 Structure Formation
PIXIE will determine the average pressure and temperature of free electrons in the post-recombination universe, providing critical input for numerical simulations of galaxy formation. Current simulations [87, 88, 89] employ a range of parameters for feedback, which couples the gravitational physics of large-scale structure to the non-linear hydrodynamics on smaller (galaxy) scales. Energy injected into circumgalactic and intracluster gas from active galactic nuclei and supernovae plays a critical but poorly-understood role in galaxy formation. Feedback processes shut off star formation in galaxies, heating and expelling gas into the intergalactic medium. The electrons in this gas distort the CMB spectrum via the thermal Sunyaev-Zel’dovich (tSZ) effect. The temperature of this gas ($10^{5}-10^{6}$ K) is sufficiently high to generate non-zero relativistic corrections to the tSZ signal [51, 87]. Assuming weak 10% priors on low-frequency foregrounds (§$\S$2.2), the baseline PIXIE mission constrains electron pressure $\sigma(y)=1.6\times 10^{-8}$ and the electron temperature $\sigma(kT_{e})=0.17$ keV, providing a new window on structure formation. Figure 10 compares the PIXIE constraints to the range of assumptions employed in simulations. Measurements of monopole $y$ quantify the amount of feedback energy injected into the ionized gas, central to refining simulations. More precise simulations, in turn, reduce uncertainty in cosmological parameter estimates from large-scale structure. Cosmological probes dependent on baryonic feedback include weak gravitational lensing and the observable-mass relation of galaxy clusters. Questions such as concordance between the amplitude of fluctuations in the early and late universe, as well as the neutrino mass sum, are hindered by baryonic feedback uncertainty. Measurements of tSZ also constrain primordial non-Gaussianity [90] with complementarity to other probes, but at lower sensitivity. For a fuller discussion, see [91].
@@ 3.5 Star Formation History
Figure 11 shows SFRD recovered from simulations using the baseline PIXIE sensitivity. PIXIE will focus on CO, [CII] and [NII] emission, which originate from dense regions of the ISM and so are good tracers of both star formation [94, 95] and metallicity [96]. PIXIE’s numerous narrow channels facilitate clean measurement across redshift [97], while the broad frequency coverage enables both [CII] and CO measurements at identical redshifts to provide independent tests of the CO vs [CII] luminosity-star formation rate relations.
@@ 3.6 Cosmic Infrared Background
PIXIE will also map the CIB distribution across the sky. The dipole anisotropy in the cosmic microwave background is attributed to the motion of the Solar System with respect to the rest frame of the surface of last scattering at redshift $z\sim 1100$. The observed homogeneity of the universe on scales larger than $\sim$100 Mpc implies that the integrated far-IR emission from galaxies at $z<6$ should show a comparable kinematic dipole, with amplitude $\Delta I/I=1.2\times 10^{-3}$ in the same direction $(l,b)=(264.021^{\circ}48.253^{\circ}$) as the CMB dipole [101]. Assuming that the CIB and CMB spectral shapes are independent of direction on the sky, a pixel-by-pixel fit of the CIB amplitude, CMB signal amplitudes, and foreground amplitudes/spectral parameters can determine the CIB dipole amplitude to 5% and direction to 4 uncertainty, where the uncertainties are dominated by confusion with the Galactic and zodiacal foregrounds.
@@ 3.7 Interstellar Medium
Interstellar dust plays critical roles in the chemical evolution of galaxies [102, 103], the processing of optical and ultraviolet light [104, 105, 106] into the infrared [107, 108, 109, 110, 111], the thermodynamics of interstellar gas [112, 113], and the formation of stars and planets [114]. It is a primary reservoir of interstellar metals, but its composition and evolution throughout the galaxy are uncertain. The first full-sky measurements of the polarized dust spectrum from Planck and the unexpectedly-high fractional polarization of the dust emission defied predictions from pre-Planck dust models [e.g., 115]. These models posit that grains exist as two distinct populations of silicate or carbonaceous grains, and thus that either ISM processing is too slow to mix them or injection timescales are short enough to maintain pristine stardust populations of distinct composition. If so, the dust spectrum in total intensity should differ from that in polarized intensity [116]. This is not observed at the microwave frequencies accessible to Planck [117] or the still higher frequencies observed by the BLASTPol balloon [118, 119]. Multi-component dust models have now been devised to respect Planck constraints [120, 121], while an alternative one-component model ("astrodust") has also been proposed [116, 122].
@@ 4 Discussion
PIXIE uses a single cryogenic Fourier transform spectrometer to measure the Stokes $I,Q,U$ parameters at sensitivity $\sim$200 Jy sr-1 in each 2.65 diameter beam over the full sky, at frequencies from 28 GHz to 6 THz. The baseline 2-year mission spends equal time measuring spectral distortions (with an external blackbody calibrator deployed to block one beam) versus polarization (with the calibrator stowed so both beams view the sky). The resulting 10.8 months of integration in each mode provides a robust detection of the spectral distortions from the mean electron pressure and temperature of the universe, and provides 95% confidence limits to B-mode polarization to detect the signal from most single-field inflationary models. The fraction of time spent in each observing mode can be changed at any point throughout the mission. PIXIE carries no expendable cryogens; a longer mission dedicated to polarization would improve the polarization limits to $r<0.9\times 10^{-3}$ to test multi-field inflation models. Conversely, a longer mission dedicated primarily to spectral distortions would lower the current 95% confidence upper limit on the chemical potential to $|\mu|<10^{-7}$. The resulting factor of 1000 improvement over the COBE/FIRAS limits opens an enormous discovery space for processes including dark matter, primordial black holes, and inflation. Appendix B summarizes instrumental improvements compared to FIRAS.
@@
The choice of a single FTS is dictated by the \$350 million USD mission budget for NASA’s MIDEX program. A larger budget would allow a more ambitious instrument architecture. This could be accomplished within NASA’s \$1B Probe line, or ESA’s L-class mission. Such a “Super-PIXIE” would employ 3 (or more) cryogenic FTS modules, each using the basic PIXIE design, but with channel width and optical passband (hence photon NEP) tuned to optimize the combined sensitivity to selected CMB signals. Table 4 shows one possible configuration. A low-frequency module would employ 5 GHz channel width over a passband 15–150 GHz to characterize the low-frequency foregrounds while avoiding photon noise from the CMB monopole at higher frequencies. A mid-frequency module would use wider 30 GHz channels over a passband 90–600 GHz to cover the cosmological signals. A high-frequency module would use 30 GHz channel width over the passband 450–6000 GHz to characterize the high-frequency foregrounds, setting the lowest frequency at 450 GHz to minimize the photon noise contribution from the CMB monopole. The two lower-frequency modules would image the sky in 7 sky pixels to increase the system etendue. Figure 14 compares the resulting sensitivity to both PIXIE and FIRAS. Such a Super-PIXIE mission could provide a $3\sigma$ detection of the minimal chemical potential distortion from dissipation of primordial density perturbations within the $\Lambda$CDM model.
@@ Appendix B Comparison to FIRAS
  • Integration Time: The FIRAS mission ended after 10 months when the liquid helium coolant ran out. Only 1 month of this time was spent observing the external blackbody calibrator, so that the sensitivity to sky spectra was limited by the signal-to-noise ratio of the calibration data. Since the single sky horn could not directly compare the sky to the external calibrator, the resulting comparison of separate sky/calibrator data sets incurred an additional $\sqrt{2}$ noise penalty. PIXIE has no expendable cryogens and can operate for extended periods. The baseline mission devotes 10.8 months to measurements of spectral distortions, during which PIXIE’s 2-beam design directly acquires the sky/calibrator difference spectrum.
@@ Appendix C Mission Noise Curves
The tables below contain the center frequencies, sensitivity, and median noise level for a map pixel for the baseline 2-year PIXIE mission. The baseline 2-year mission assumes that 45% of observing time is spent with the calibrator deployed (sensitive to spectral distortions and polarization) and 45% with the calibrator stowed (sensitive to polarization only). Entries for polarization noise (Table LABEL:pol_table) include the contribution from both modes. Noise levels assume 30% of the observing time in each mode is spent with narrow synthesized channel width and 70% with broad channel width (§$\S$2.1). The channel widths depend on the MTM mirror stroke and can be changed (within hardware limits) throughout the mission.

### v2

@@ 3.3 Particle Physics
The neutrino is the only Standard Model particle whose mass is unknown. Measurements of E-mode polarization enable detection of the neutrino mass. Neutrinos affect the growth of structure: massive neutrinos cannot cluster on physical scales smaller than their free-streaming length, suppressing the growth of structure on these scales. Gravitational lensing from CMB and optical surveys determines the amplitude of current structure in our universe, but determining the growth of structure requires knowledge of the primordial amplitude as well. Primary CMB anisotropy only determines the primordial amplitude within $\sim$1uncertainty due to scattering at reionization. PIXIE’s nearly cosmic-variance limited measurement of the reionization optical depth (§$\S$2.2) would remove this limiting uncertainty to enable a determination of neutrino mass to 4 standard deviation precision [83].
Dark matter (DM) provides a more exotic test of particle physics. The long-favored weakly interacting massive particles (WIMP) scenario is seeing increased pressure from both direct detection experiments and tests for supersymmetry in particle colliders, suggesting instead more exotic scenarios such as axionic DM [84, 85] or PBHs [86]. PIXIE’s spectral distortion limits constrain electromagnetic interactions between dark matter and standard-model particles from DM decay annihilation and scattering [87]. Spectral distortions from DM decay depend on the lifetime and abundance of the (excited) DM state. The temperature/velocity- dependence of the DM annihilation cross-section strongly affects annihilation signals, with spectral distortions being particularly valuable for p-wave annihilation [88]. Since the dark matter number density scales as 1/mass, spectral distortions from DM annihilation or scattering are sensitive to low masses. PIXIE 95% confidence limits $|\mu|<5\times 10^{-7}$ from the baseline mission restrict Majorana annihilations to mass $m_{\chi}>{\sim}300$ keV [89].
@@ 3.4 Structure Formation
PIXIE will determine the average pressure and temperature of free electrons in the post-recombination universe, providing critical input for numerical simulations of galaxy formation. Current simulations [90, 91, 92] employ a range of parameters for feedback, which couples the gravitational physics of large-scale structure to the non-linear hydrodynamics on smaller (galaxy) scales. Energy injected into circumgalactic and intracluster gas from active galactic nuclei and supernovae plays a critical but poorly-understood role in galaxy formation. Feedback processes shut off star formation in galaxies, heating and expelling gas into the intergalactic medium. The electrons in this gas distort the CMB spectrum via the thermal Sunyaev-Zel’dovich (tSZ) effect. The temperature of this gas ($10^{5}-10^{6}$ K) is sufficiently high to generate non-zero relativistic corrections to the tSZ signal [52, 90]. Assuming weak 10% priors on low-frequency foregrounds (§$\S$2.2), the baseline PIXIE mission constrains electron pressure $\sigma(y)=1.6\times 10^{-8}$ and the electron temperature $\sigma(kT_{e})=0.17$ keV, providing a new window on structure formation.
Figure 10 compares the PIXIE constraints to the range of assumptions employed in simulations. We adopt the baseline PIXIE mission noise curves (Appendix C) with a total of 7.6 months integration time on the cleanest 70% of the sky. In addition to the $y$-distortion and its relativistic correction, the Fisher forecast includes six astrophysical foregrounds with 11 free parameters as described in §$\S$ 2.2. Additionally, $10\%$ priors are assumed for both the synchrotron and free-free emission, and no curvature parameter for the synchrotron emission is included in this setup. The resulting limits differ from [37] in that both an updated noise curve and a different baseline integration time have been used.
Measurements of monopole $y$ quantify the amount of feedback energy injected into the ionized gas, central to refining simulations. As can be seen in Fig. 10, current models of astrophysical feedback depend on highly uncertain parameters, yielding a wide range of predictions in the $y-T_{e}$ plane. More precise simulations reduce uncertainty in cosmological parameter estimates from large-scale structure. Cosmological probes dependent on baryonic feedback include weak gravitational lensing and the observable-mass relation of galaxy clusters. Questions such as concordance between the amplitude of fluctuations in the early and late universe, as well as the neutrino mass sum, are hindered by baryonic feedback uncertainty. Measurements of tSZ also constrain primordial non-Gaussianity [93] with complementarity to other probes, but at lower sensitivity. For a fuller discussion, see [94].
@@ 3.5 Star Formation History
Figure 11 shows SFRD recovered from simulations using the baseline PIXIE sensitivity. for the fiducial SFRD model [96]. PIXIE will measure both [CII] and CO emission. We assume the [CII] luminosity model from [97]
$L_{\rm CII}(M,z)=\left(\frac{M}{M_{1}}\right)^{\beta}\exp(-N_{1}/M)\left(\frac{(1+z)^{2.7}}{1+[(1+z)/2.9)]^{5.6}}\right)^{\alpha}$ (3.1)
such that the [CII] luminosity is related to the SFR by a power law, $L\propto{\rm SFR}^{\alpha}$. The parameters $\beta$, $N_{1}$, and $M_{1}$ are set at all redshifts using priors from [97]. We set $\alpha$ at each redshift in a different manner. We use the Santa Cruz semi-analytic model (SAM) [98, 95, 99], which predicts both [CII] luminosities and SFRs for galaxies, to fit $\alpha$ at each redshift using [CII] intensities and SFRDs from the SAM. For the SFRD forecast from PIXIE’s CO(5-4) intensity measurement, we assume the Yang et al. (2019) CO(5-4) luminosity model, which is a fit to the Santa Cruz SAM. In this case we assume a linear relationship between the CO(5-4) luminosity and the SFR, such that the CO(5-4) intensity is proportional to the SFRD.
PIXIE will focus on CO, [CII] and [NII] emission, which originate from dense regions of the ISM and so are good tracers of both star formation [100, 101] and metallicity [102]. PIXIE’s numerous narrow channels facilitate clean measurement across redshift [103], while the broad frequency coverage enables both [CII] and CO measurements at identical redshifts to provide independent tests of the CO vs [CII] luminosity-star formation rate relations.
@@ 3.6 Cosmic Infrared Background
PIXIE will also map the CIB distribution across the sky. The dipole anisotropy in the cosmic microwave background is usually attributed to the motion of the Solar System with respect to the rest frame of the surface of last scattering at redshift $z\sim 1100$. The observed homogeneity of the universe on scales larger than $\sim$100 Mpc implies that the integrated far-IR emission from galaxies at $z<6$ should show a comparable kinematic dipole, with amplitude $\Delta I/I=1.2\times 10^{-3}$ in the same direction $(l,b)=(264.021^{\circ},48.253^{\circ}$) as the CMB dipole [107]. Since the CIB represents the integrated emission from all galaxies, its dipole does not suffer from potential bias from survey completeness. A statistically significant difference in either amplitude or direction between the CMB and CIB dipoles would provide evidence for a non-kinematic component for the CMB dipole resulting from pre-inflationary physics [108, 109].
@@
To quantify the PIXIE sensitivity to the CIB dipole, we generate a sky model with the nominal CMB, CIB, and Galactic foregrounds, including the dominant foregrounds from Galactic dust and zodiacal emission. We then add simulated instrument noise and perform a pixel-by-pixel fit of the CIB amplitude, CMB amplitude, and foreground amplitudes/spectral parameters (§$\S$2.2) for 5000 realizations of the noise. We assume that the CIB spectral energy distribution is independent of direction on the sky, fitting a global CIB spectrum with amplitude fitted individually in each pixel. Foreground parameters (amplitude, temperatures, spectral indices) are fit on a pixel-by-pixel basis. The resulting distribution for the CIB in each pixel recovers the CIB dipole amplitude to 5% and direction to 4 uncertainty, where the uncertainties are dominated by confusion with the Galactic and zodiacal foregrounds.
@@ 3.7 Interstellar Medium
Interstellar dust plays critical roles in the chemical evolution of galaxies [110, 111], the processing of optical and ultraviolet light [112, 113, 114] into the infrared [115, 116, 117, 118, 119], the thermodynamics of interstellar gas [120, 121], and the formation of stars and planets [122]. It is a primary reservoir of interstellar metals, but its composition and evolution throughout the galaxy are uncertain. The first full-sky measurements of the polarized dust spectrum from Planck and the unexpectedly-high fractional polarization of the dust emission defied predictions from pre-Planck dust models [e.g., 123]. These models posit that grains exist as two distinct populations of silicate or carbonaceous grains, and thus that either ISM processing is too slow to mix them or injection timescales are short enough to maintain pristine stardust populations of distinct composition. If so, the dust spectrum in total intensity should differ from that in polarized intensity [124]. This is not observed at the microwave frequencies accessible to Planck [60] or the still higher frequencies observed by the BLASTPol balloon [125, 126]. Multi-component dust models have now been devised to respect Planck constraints [127, 128], while an alternative one-component model ("astrodust") has also been proposed [124, 129].
@@
Table 4: Line Emission Signal To Noise Ratio
Line Median SNRa
CO J=1-0 0.1
CO J=2-1 2.2
CO J=3-2 10
CO J=4-3 26
CO J=5-4 27
CO J=6-5 26
CO J=7-6 11
C+ 158 um 5040
N+ 205 um 450
aMedian per-pixel signal to noise ratio at $|b|>30^{\circ}$
PIXIE will also map line emission from Galactic molecular, atomic, and ionic species. Although the moderate spectral resolution prevents velocity-resolved measurements in the Galactic plane, PIXIE will map the velocity-integrated line emission over the full sky. We use low-SNR maps of the CO lines from Planck and the singly-ionized carbon and nitrogen lines from FIRAS to estimate the PIXIE line sensitivity at latitudes $|b|>20^{\circ}$. Table 4 lists the median per-pixel signal to noise ratio for these lines.
@@ 4 Discussion
PIXIE uses a single cryogenic Fourier transform spectrometer to measure the Stokes $I,Q,U$ parameters at sensitivity $\sim$200 Jy sr-1 in each 2.65 diameter beam over the full sky, at frequencies from 28 GHz to 6 THz. The baseline 2-year mission spends equal time measuring spectral distortions (with an external blackbody calibrator deployed to block one beam) versus polarization (with the calibrator stowed so both beams view the sky). The resulting 7.6 months of integration in each mode over the cleanest 70% of the sky provides a robust detection of the spectral distortions from the mean electron pressure and temperature of the universe, and provides 95% confidence limits to B-mode polarization to detect the signal predicted from most single-field inflationary models. The fraction of time spent in each observing mode can be changed at any point throughout the mission. PIXIE carries no expendable cryogens; a longer mission dedicated to polarization would improve the polarization limits to $r<0.9\times 10^{-3}$ to test multi-field inflation models. Conversely, a longer mission dedicated primarily to spectral distortions would lower the current 95% confidence upper limit on the chemical potential to $|\mu|<10^{-7}$. The resulting factor of 1000 improvement over the COBE/FIRAS limits opens an enormous discovery space for processes including dark matter, primordial black holes, and inflation. Appendix B summarizes instrumental improvements compared to FIRAS.
The choice of a single FTS is dictated by the \$300 million (2022 USD) mission budget for NASA’s MIDEX program.333 https://explorers.larc.nasa.gov/2021APMIDEX/MIDEX/index.html A larger budget would allow a more ambitious instrument architecture. This could be accomplished within NASA’s \$1B Astrophysics Probe line444 https://explorers.larc.nasa.gov/2023APPROBE/announcements.html or ESA’s L-class mission. Such a “Super-PIXIE” would employ 3 (or more) cryogenic FTS modules, each using the basic PIXIE design, but with channel width and optical passband (hence photon NEP) tuned to optimize the combined sensitivity to selected CMB signals. Table 5 shows one possible configuration. A low-frequency module would employ 5 GHz channel width over a passband 15–150 GHz to characterize the low-frequency foregrounds while avoiding photon noise from the CMB monopole at higher frequencies. A mid-frequency module would use wider 30 GHz channels over a passband 90–600 GHz to cover the cosmological signals. A high-frequency module would use 30 GHz channel width over the passband 450–6000 GHz to characterize the high-frequency foregrounds, setting the lowest frequency at 450 GHz to minimize the photon noise contribution from the CMB monopole. The two lower-frequency modules would image the sky in 7 sky pixels to increase the system etendue. Figure 14 compares the resulting sensitivity to both PIXIE and FIRAS. Such a Super-PIXIE mission could provide a $3\sigma$ detection of the minimal chemical potential distortion from dissipation of primordial density perturbations within the $\Lambda$CDM model.
@@ Appendix B Comparison to FIRAS
  • Integration Time: The FIRAS mission ended after 10 months when the liquid helium coolant ran out. Only 1 month of this time was spent observing the external blackbody calibrator, so that the sensitivity to sky spectra was limited by the signal-to-noise ratio of the calibration data. Since the single sky horn could not directly compare the sky to the external calibrator, the resulting comparison of separate sky/calibrator data sets incurred an additional $\sqrt{2}$ noise penalty. PIXIE has no expendable cryogens and can operate for extended periods. The baseline mission devotes 10.8 months to measurements of spectral distortions over the full sky, during which PIXIE’s 2-beam design directly acquires the sky/calibrator difference spectrum.
@@ Appendix C Mission Noise Curves
The tables below contain the center frequencies, sensitivity, and median noise level for a map pixel for the baseline 2-year PIXIE mission. The baseline 2-year mission assumes that 45% of observing time is spent with the calibrator deployed (sensitive to spectral distortions and polarization) and 45% with the calibrator stowed (sensitive to polarization only). Entries for polarization noise (Table LABEL:pol_table) include the polarization contribution from both modes. Noise levels assume 30% of the observing time in each mode is spent with narrow synthesized channel width (MTM long-stroke mode) and 70% with broad channel width (MTM short-stroke mode). See §$\S$2.1 for details. The channel widths depend on the MTM mirror stroke and can be changed (within hardware limits) throughout the mission.